ZZ, Inc. has budgeted sales in units for the next six months as follows:

                  Budgeted Sales in Units
July                    7,700 units                             
August                  8,300 units
September               5,400 units
October                10,200 units
November               ?????? units
December               11,700 units

The selling price is $18 per unit. 30% of the company's sales are cash
sales and 70% of the company's sales are made on account. The sales on
account are collected in the pattern 15% in the month of sale, 20% in
the month following sale, 55% in the second month following sale, and
the final 10% is collected in the third month following sale.

ZZ, Inc. reported a budgeted accounts receivable balance of $194,670
in its pro forma balance sheet at December 31.

Calculate the number of units budgeted to be sold in November. Do not
type the word units after your answer.
Answer #1
Accounts receivable balance 194670
Less: October sales uncollected 12852 =10200*18*70%*10%
Less: December sales uncollected 125307 =11700*18*70%*85%
November sales uncollected 56511
Percentage of November sales uncollected 65% =55%+10%
Credit sales fo November 86940 =56511/65%
Total sales for November 124200 =86940/70%
Number of units budgeted to be sold in November 6900 =124200/18
Answer #2
Budgeted accounts receivable 194670
Less: Uncollected October sales 12852 =10200*18*70%*10%
Less: Uncollected December sales 125307 =11700*18*70%*85%
Uncollected November sales 56511
Uncollected November sales percentage 65%
November credit sales 86940 =56511/65%
November Total sales 124200 =86940/70%
Number of units budgeted to be sold in November 6900 =124200/18